Given the utility function:
U = X 3/4 . Y1/4
Find out the optimal quantities of the two commodities using Lagrange method and simplified procedure, if it is given that price of X is Rs.6 and price of Y is Rs.3 and income is equal to Rs.120.
1
Expert's answer
2020-08-18T12:50:16-0400

Income = 120

Objective: Maximize U = "X^{0.75}*Y^{0.25}"

Subject to: Total Revenue (TR) = "P_X*X +P_Y*Y"

                       "120 = 6*X +3*Y"

L(λ, x, y) = "X^{0.75}*Y^{0.25} \u2013 \u03bb(6X +3Y-120)"

"\frac{∂L}{∂X}" = "0.75X^{-0.25}*Y^{0.25} - \u03bb(6) = 0"

"\frac{∂L}{∂Y}" = "0.25X^{0.75}*Y^{-0.75} \u2013 \u03bb (3) = 0"

"\frac{∂L}{∂ λ}" = -1(6X +3Y-120) = 0

Ratio of First order condition (FOC)

"\frac{∂L}{∂X} /\frac{∂L}{∂Y}" = "\frac{λ6}{λ3} = 0.75X^{-0.25}*Y^{0.25} /0.25X^{0.75}*Y^{-0.75}"

"\frac{6}{3}" = "\frac{0.75}{0.25}X^{-0.25-0.75}*Y^{0.25- -0.75}"

2 ="3X^{-1}*Y^{1}"

Y = "\frac{2}{3}X"

Solving for the equations:

6X +3Y= 120

"Y = \frac{2}{3}X"

"6X +3*\frac{2}{3}X = 120"

6X +2X = 120

8X = 120

X = "\frac{120}{8} = 15"

Solving for y

Y = "\frac{2}{3}X"

Y = "\frac{2}{3}*15"

Y = 10

Solving for Maximum Utility

U = "X^{0.75}*Y^{0.25}"

X = 15

Y = 10

U = "15^{0.75}*10^{0.25}"

U = "7.62199*1.77828"

U = 13.554